# envname-P envname#1

### v2

@@ Abstract
Prevailing code generation benchmarks, such as HumanEval+ and MBPP+, primarily evaluate large language models (LLMs) with pass@k on functional correctness using well-formed inputs. However, they ignore a crucial aspect of real-world software: adherence to contracts—the preconditions and validity constraints that dictate how ill-formed inputs must be rejected. This critical oversight means that existing benchmarks fail to measure, and models consequently fail to generate, truly robust and reliable code snippets. We introduce PACT, a program assessment and contract-adherence evaluation framework, to bridge this gap. PACT is the first framework designed to systematically evaluate and enhance contract-adherence in LLM-generated code snippets alongside functional correctness. PACT’s contributions are threefold: First, it provides a comprehensive test-suite corpus focused on contract violations, extending HumanEval+ and MBPP+. Second, it enables a systematic analysis of code generation under varied prompting conditions. This analysis demonstrates that augmenting prompts with contract-violating test cases significantly enhance a model’s ability to respect contracts compared to using contract description alone. Finally, it introduces novel metrics to rigorously quantify contract adherence in both test generation and code generation. By revealing critical errors that conventional benchmarks overlook, PACT provides the rigorous and interpretable metrics to evaluate the robustness of LLM-generated code snippets in both functionality and contract-adherence. Our code and data are available at https://github.com/suhanmen/PACT.
@@ 1 Introduction
Test case generation is particularly crucial since it validates correctness, reveals corner cases, and supports automated evaluation. Traditional approaches such as random testing (Cha et al., 2015), symbolic execution (King, 1976; Yoon & Cha, 2024; Cadar et al., 2008), and search-based software testing (Harman & McMinn, 2010; Formica et al., 2024; Burnim & Sen, 2008) have laid the groundwork. Then, various large language model (LLM)-driven methods have recently been investigated, including zero-shot prompting (Chen et al., 2021a; Jaremko et al., 2025), coverage-guided test generation (Ryan et al., 2024; Lemieux et al., 2023; Sapozhnikov et al., 2024), mutation-informed generation (Dakhel et al., 2024; Cajica et al., 2021), slicing-based decomposition (Wang et al., 2024), and satisfiability modulo theories (SMT) solver-based test generation (Peleska et al., 2011; Cadar et al., 2008).
Yet current test case generation approaches (Srivastava & Payer, 2021; Yoon & Cha, 2024) and the resulting benchmarks (Chen et al., 2021a) still judge the quality of both test cases and programs exclusively through pass@k (Chen et al., 2021a). In practice, real-world software is governed by contracts (Meyer, 1992)—the preconditions and input validation rules that restrict valid inputs and define expected behavior for ill-formed ones. As a black-box metric, pass@k measures correctness only via input-output correspondence on well-formed inputs, which are test cases that have already been filtered to comply with these contracts. Consequently, evaluating code snippets solely by well-formed input-output correspondence without considering contracts leads to an inaccurate assessment, as it overlooks whether the generated code snippets enforce pre-conditions and input validation checks explicitly stated or implicitly included in the specification (Liu et al., 2023).
Typically, most coding competitions such as International Collegiate Programming Contest  [^1] ICPC website: https://icpc.global/ (ICPC) or International Olympiad in Informatics  [^2] IOI website: https://ioinformatics.org/ provide specifications that state the constraints to consider when implementing the function for the task. This description includes not only the functional goal—such as computing a sum or sorting a list—but also a set of explicit contracts, such as input bounds, type restrictions, and required error-handling behaviors. These contracts often phrased in natural language or implied via examples, act as contracts that describe the valid input space and define the expected behavior under edge or invalid conditions. Figure 1 provides a clear illustration, showing how a simple functional task description contains hidden contracts that are often overlooked. For instance, a function designed to find duplicate values in a list of numbers, has_dup(nums), has an implicit contract that the input must be a list containing only integers. A correct implementation should therefore include specific assertions—such as assert isinstance(nums, list) and assert all(isinstance(x, int) for x in nums)—to verify the input’s type before proceeding with the functional logic. If such checks are omitted, the program may appear correct on well-formed inputs like [1, 2, 3]. However, it will fail to reject ill-formed inputs like ’a’ or [’2’, 2], thus failing to reflect the intended specification. Ignoring these contracts in code generation poses serious reliability and safety issues. For instance, a module might silently accept an invalid input, such as a negative number where a positive one is expected. By doing so, it can propagate a hidden fault by passing a nonsensical result to another part of the system. This chain reaction often leads to logical errors that are extremely difficult to trace back to their original source.
Figure 1: PACT’s contract-violating test uncovers an implicit constraint that conventional functional tests miss, proving the need for contract-aware evaluation.
Therefore, contracts are not optional safeguards but integral parts of the specification that define the boundary of valid behavior. Without them, even code snippets that appear functionally correct cannot be trusted to operate safely and reliably. Empirical studies reveal that LLM-generated code snippets frequently overlook these contracts and often fails to guard against ill-formed inputs that violate them. As a result, a generated code may incorrectly pass tests that it should fail, creating the illusion of correctness under default test cases. This shows that many seemingly correct solutions in fact ignore contractual requirements, underscoring that contract adherence is a fundamental necessity for trustworthy code generation.
We introduce PACT, a program assessment and contract-adherence evaluation framework, to bridge this gap. The main focus of PACT is twofold: first, to construct high-quality contract-violating test cases, and second, to use them for a systematic analysis of contract-aware code generation. For test case generation, our approach leverages an SMT solver to systematically explore diverse combinations of contract violations, moving beyond simplistic, single-violation tests.
For example, for a function with the contracts assert isinstance(r, float), assert r > 0, and assert h > 0, our method can generate an input such as r = -2.5 and h = 5. This test case is guaranteed to satisfy the type contract for r and the positivity contract for h, while precisely violating only the positivity contract for r. By pruning logically inconsistent combinations in advance, the solver ensures that only feasible and semantically valid test cases are generated. For our analysis of code generation, we then use these precisely targeted test cases to evaluate how different prompting strategies guide an LLM to produce implementations that properly enforce the intended contracts.
In summary, current benchmarks overlook contract violations, thereby inflating the perception of correctness. Addressing this limitation, we propose the PACT framework. PACT extends HumanEval+ and MBPP+ to generate test cases focused on contract violations and introduces novel metrics to quantify contract adherence.
@@ 2 Related work
2.1 Automated Test and Code Generation
Automated test and code generation have deep roots in traditional software engineering. Early test case generation was typically categorized into black-box methods like random and mutational fuzzing (Godefroid et al., 2005; Claessen & Hughes, 2000), white-box techniques such as symbolic execution (He et al., 2021; Cha et al., 2022; Godefroid et al., 2008), and grey-box approaches like coverage-guided fuzzing (Choi et al., 2019; Stephens et al., 2016; She et al., 2024; Qian et al., 2022). Similarly, early code generation relied on methods like probabilistic grammar-based frameworks (Bielik et al., 2016) and specialized language models (Feng et al., 2020). The advent of the Transformer architecture marked a paradigm shift, establishing LLMs as the dominant approach in both fields (Chen et al., 2021b). The development of modern Code LLMs now typically involves pre-training on code corpora, followed by instruction tuning (Ouyang et al., 2022) and refinement through more advanced techniques like reinforcement learning with execution feedback (Gehring et al., 2025). However, despite these advancements, the primary goal has remained the enhancement of functional correctness, typically measured by pass@k on benchmarks with well-formed inputs.
2.2 Contracts
The Design by Contract (DbC) paradigm (Meyer, 1992) argues that software reliability depends on explicitly stated preconditions, postconditions, and invariants, with run-time assertions as the enforcement mechanism. However, mainstream automated testing and existing test-generation benchmarks (Wang et al., 2025; Jain et al., 2025) often overlook latent defects from missing contract checks. While related work on failure-inducing test cases (Zhang et al., 2024; Zhong et al., 2025) is effective at causing general exceptions, this is not the same as verifying adherence to the precise semantics of a given contract. For instance, for a function that requires a list of positive numbers, a generic failure-inducing test might use None to cause a TypeError, but this does not verify the specific rule that all numbers must be positive. In contrast, a contract-aware test like [10, 20, -5] is designed to be caught precisely by an assertion checking for positive values. This distinction highlights the need for systematic methods that can precisely target formal contract specifications rather than just triggering arbitrary errors.
2.3 SMT Solver
SMT (Satisfiability Modulo Theories) solvers are powerful engines for determining the satisfiability of complex logical formulas across various theories, such as arithmetic and strings (Barrett & Tinelli, 2018). They typically interface using SMT-LIB, a standardized formal language, with Z3 being a widely adopted implementation in automated testing (de Moura & Bjørner, 2008). Traditionally, techniques like symbolic execution have employed SMT solvers to find inputs that satisfy path constraints, focusing on functional coverage over well-formed inputs. In this paradigm, constraints are used as admission checks to filter out invalid data rather than as explicit targets for evaluation. In contrast, our framework leverages SMT solvers to systematically generate ill-formed inputs that precisely violate formalized contracts, allowing us to rigorously test for robustness. A detailed example of how we formulate contracts is provided in Appendix D.
3 The Need for Contract-Aware Evaluation
3.1 Why Contracts Matter: Blind Spots in Current Benchmarks
Recent studies on code and test case generation (Korraprolu et al., 2025; Sung et al., 2025; Li et al., 2022) have dominantly relied on pass@k to evaluate functional correctness. While effective, these metrics, along with approaches that target failure-inducing inputs (Peng et al., 2018), share a common limitation: they primarily operate within the bounds of a program’s legal input space Dyck et al. (2023). This practice has led to a critical blind spot in popular benchmarks such as HumanEval+ and MBPP+. By design, these benchmarks explicitly filter out and discard any test case that violates a program’s pre-conditions (Austin et al., 2021). Consequently, the evaluation process certifies that a program behaves correctly on well-formed inputs, but it fails entirely to assess the program’s robustness against ill-formed ones. This results in an incomplete and often inflated assessment of code quality, praising solutions that may be superficially correct but are fundamentally fragile.
This is where the software contracts become essential. As explained in Section 1, contracts are the rules that define the boundary of valid behavior. They are not optional safeguards but a core component of trustworthy software that specifies how a program must identify and reject invalid data. By ignoring contract adherence, existing benchmarks overlook a crucial dimension of software reliability.
3.2 Contract-Based Test Paradigm
Merely enlarging a single pool of valid inputs cannot reveal whether a model understands the boundary of the specification it needs to satisfy. The main focus, therefore, lies in constructing contract-violating test cases, which systematically explore the extent to which models enforce contract rules. A contract-violating test case is an input that violates one or more predicates from a set of contracts while satisfying the remaining specification. The reference implementation is augmented with run-time assertions for every predicate in the contract set, and a candidate program passes such a test only when it raises an error consistent with this augmented reference. Introducing contract-violating inputs uncovers false negatives that purely functional tests overlook and provides a rigorous measure of whether a program properly enforces contractual rules.
3.3 Task Setup
Each benchmark task consists of a natural language description, a set of contracts, and a functional implementation. We use HumanEval+ and MBPP+, where contract predicates are provided as assertion literals outside the prompt and reference code. Our task concerns contract-violating test generation and contract-aware code generation. For test generation, we automatically construct a compact set of contract-violating test cases that target specific contract rules and remain feasible with respect to the remaining rules for each benchmark task. These tests are used for evaluating whether LLM-generated code appropriately follows contracts. For contract-aware code generation, we generate code under two prompt conditions, where the first condition is a contract specification (CS) prompt, which includes the functional description and a natural language paraphrase of the contracts. The second condition is an example-augmented specification (EAS) prompt, which is augmented with the contract-violating test cases.
Figure 2: Running example of PACT with an Example-Augmented Specification (EAS) prompt, which integrates Contract Specification (CS) and contract-violating test cases (CVTs) to enforce contract-aware code generation.
4 Methodology
We propose PACT, a program assessment and contract-adherence evaluation framework designed to systematically evaluate and enhance the ability of LLMs to generate contract-compliant code. This framework consists of two main stages. First, we generate contract-violating test cases to rigorously assess whether LLM-generated code snippets enforce both functional specifications and explicit contracts. Second is the systematic analysis of code generation, where we use these test cases under different prompting conditions to evaluate a model’s contract awareness in detail. Unlike prior approaches that rely solely on functionality-based evaluation, PACT extends the evaluation paradigm with contract-violation tests, enabling a more precise and reliable analysis of contract adherence.
4.1 Contract-Violating Test-Case Generation
Direct LLM-based generation is inadequate for contract-aware test case construction. It must consider all subsets of the contract set when producing violating inputs. If a task has $n$ contracts, the number of non-empty violation combinations is $2^{n}-1$. Direct prompting often misses required combinations and yields contradictory inputs that violate unintended contracts and fails to ensure feasibility under the specification. LLMs also lack a built-in mechanism to verify that exactly the targeted contracts are violated while all others hold. We design an SMT-based approach to make these checks accurate and efficient. Our generation procedure is a two-step pipeline for contract-aware test case generation. First, an inference model translates natural language contracts into an Algebraic Data Type (ADT) program. This ADT format is the critical component: it provides a rigid, formal schema for the complex and often nested constraints found in contracts. This ensures that the subsequently generated rules for an SMT solver are syntactically valid and semantically precise. Second, given the ADT, the SMT solver constructs contract violation test cases (CVTs) by breaking one or more specified contracts while ensuring that all remaining contracts are satisfied. The solver validates satisfiability for each candidate combination and extracts concrete models—specific assignments of values that satisfy the constraints— to instantiate inputs for every valid case. We provide a running example of this procedure in Appendix D.
4.2 Contract Metrics for Test Case
Unlike standard test cases for functional correctness, test cases for contracts should be the negative samples, violating the contracts and triggering the corresponding assertions. As the concept of contract-violating test cases is different from the standard, we design metrics to analyze whether the generated test cases appropriately correspond to input contracts.

### v3

# ContractEval: A Benchmark for Evaluating Contract-Satisfying Assertions in Code Generation

@@ Abstract
Current code generation benchmarks measure functional correctness on well-formed inputs, as test cases are curated to satisfy input preconditions. This leaves a gap: generated programs may appear correct but fail to satisfy contracts—assertion-level validity constraints for rejecting ill-formed inputs. We introduce ContractEval, a benchmark for evaluating contract-satisfying assertions in code generation, i.e., whether code rejects contract-violating inputs by triggering intended assertions. Built on HumanEval+ and MBPP+, ContractEval augments each task with contract-violation tests derived from reference assertions. We synthesize these via a neuro-symbolic pipeline: an LLM converts assertion clauses into constraints, and an SMT solver enumerates satisfiable violation combinations to generate inputs that violate selected clauses while satisfying the rest. Across five code LLMs, standard prompting yields 0% contract satisfaction on contract-violating test cases. Adding a small number of contract-violation examples increases contract satisfaction to 49–53% and retains 92% of the original pass@1. Our code is available at https://github.com/suhanmen/ContractEval.
[…]
envname-P envname#1
ContractEval: A Benchmark for Evaluating Contract-Satisfying Assertions in Code Generation
1Soohan Limthanks: Equal contribution.  ,   1Joonghyuk Hahn11footnotemark: 1  ,   2Hyunwoo Park  ,   2Sang-Ki Ko  ,   1Yo-Sub Hanthanks: Corresponding author. 1Yonsei University, Seoul, Republic of Korea  2University of Seoul, Seoul, Republic of Korea {aness1219,greghahn,emmous}@yonsei.ac.kr , {hwpark03,sangkiko}@uos.ac.kr
@@ 1 Introduction
Current code generation benchmarks (Chen et al., 2021; Austin et al., 2021; Liu et al., 2023) primarily measure functional correctness on well-formed inputs, typically through pass@k (Kulal et al., 2019; Chen et al., 2021). In this setting, test cases are curated to satisfy (often implicit) input conditions, and generated programs are judged by input-output correspondence on these pre-filtered inputs. This evaluation is appropriate for verifying functional behavior within the intended input domain, but it leaves a critical blind spot: a program can look correct under pass@k while failing to reject ill-formed inputs that violate validity constraints. Considering real-world software, failing to rule out such inputs and accepting invalid input can silently propagate erroneous states and undermine reliability assumptions.
A natural way to formalize this boundary is through contracts (Meyer, 1992). In this work, we focus on assertion-level contracts—preconditions and input validity constraints that are realized as explicit assertion checks in a code (e.g., type, shape, range constraints). Importantly, many widely used benchmarks implicitly rely on such contracts during dataset constructions: ill-formed inputs are filtered out, leaving only well-formed tests for evaluation (Liu et al., 2023). As a result, it is possible for a generated solution to pass functional tests while still accepting inputs that a developer expects the implementation to reject. Figure 1 illustrates this mismatch.
We introduce ContractEval, a benchmark for evaluating whether generated programs implement contract checks and reject ill-formed inputs. ContractEval is built on HumanEval+ and MBPP+ (Liu et al., 2023) by augmenting each task with contract-violation (ill-formed) tests derived from reference assertions. We define rejection as raising AssertionError. Because available reference assertions are not guaranteed to be complete, ContractEval includes a contract-quality filtering step to remove misaligned contracts and evaluates contract satisfaction with respect to the aligned assertions,
A key technical challenge is generating contract-violation tests that target specific assertions. When multiple assertions interact, violating a chosen subset while satisfying the rest requires a sophisticated procedure to reflect dependencies among contract assertions. ContractEval addresses this via a neuro-symbolic pipeline that combines an LLM and an SMT solver (Barrett and Tinelli, 2018). An LLM converts assertions into SMT-compatible constraints, and an SMT solver enumerates satisfiable violation combinations to generate concrete inputs that violate selected assertions while satisfying others. This yields contract-targeted counterexamples that support fine-grained contract evaluation.
Using ContractEval, we find that standard prompting frequently produces programs that pass well-formed tests but fail to reject contract-violating inputs. We further show that injecting a small number of contract-violation examples into the prompt substantially improves contract satisfaction while largely preserving functional correctness measured by pass@1.
Figure 1: ContractEval provides test cases and metrics to evaluate whether the generated code satisfies contracts.
@@ 2 Related work
2.1 Code Generation Evaluation
Most code generation benchmarks evaluate models by executing generated programs on unit tests and reporting pass@k (Wang et al., 2025; Jain et al., 2025), which captures functional input-output correctness on well-formed inputs (Chen et al., 2021; Hendrycks et al., 2021; Li et al., 2022). This evaluation standard is effective for verifying behavior within the intended input domain, but it typically does not measure whether a generated program reject ill-formed inputs that violate implicit validity constraints. Recent benchmark extensions (Chen et al., 2023; Liu et al., 2023) improve the coverage of functional behaviors by adding stronger or more diverse valid tests, yet their evaluation suites still primarily target well-formed inputs, leaving contract satisfaction largely unmeasured.
2.2 Contract Validity with Assertions
Design by Contract (DbC) treats preconditions, postconditions, and invariants as part of the specification, and emphasizes checking them in executable conditions such as assertion (Meyer, 1992). Contracts most commonly appear as assertion-level input validity checks that represent the intended input space. Prior work on robustness-oriented testing studies failure-inducing inputs that trigger generic exceptions (Zhang et al., 2024; Zhong et al., 2025), but a crashing input does not necessarily correspond to any particular contract assertion. For instance, assume a function that requires a list of positive numbers. An input None may trigger a TypeError regardless of the intended rule, “all numbers must be positive.” In contrast, contract evaluation requires tests that intentionally violate a chosen contract while remaining feasible with respect to others. This distinction highlights the need for systematic methods that can precisely target formal contract specifications rather than just triggering arbitrary errors. ContractEval targets this assertion-level notion of contracts and evaluates whether the generated programs reject ill-formed inputs in the intended way, using reference assertions available in HumanEval+ and MBPP+ as a starting point.
2.3 SMT Solver for Test Generation
SMT (Satisfiability Modulo Theories) solvers decide the satisfiability of logical formulas under background theories (e.g., arithmetic, strings), and are widely used in program analysis and constraint-based test generation (Barrett and Tinelli, 2018; de Moura and Bjørner, 2008). Classic uses include symbolic execution and path-constraint solving to synthesize inputs that satisfy feasibility constraints (Cadar et al., 2008). ContractEval uses SMT solving in a complementary way: we convert contract assertions into solver constraints and query for test cases that violate a selected subset of the assertions while satisfying the rest. This enables contract-targeted ill-formed test generation, which is difficult to achieve reliably with LLM-only generation.
3 Observation on Contract Satisfaction
HumanEval+ and MBPP+ provide assertion-level contracts as a dedicated dataset field, but their default test suites exclude contract-violating inputs by construction. As a result, standard evaluation certifies functional behavior only on well-formed inputs and leaves contract satisfaction untested. We expose this blind spot by executing generated implementations on contract-violating tests (CVTs) and measuring whether the implementation rejects them.
We compare a standard benchmark prompt and a prompt augmented with negative examples. Figure 2 shows that the generated code rejects none of the CVTs under both prompts, indicating that negative examples do not elicit reliable contract checks. We then evaluate a contract-aware prompt that states the intended input constraints, and the same models reject substantially more CVTs under this prompt. This suggests that commonly used benchmark prompts often fail to generate programs that follow the provided input constraints in practice, unless they are surfaced explicitly.
Figure 2: Evaluates contract satisfaction (contract sat.) of generated code. Standard and negative-example prompts yields 0% while a contract prompt achieves 50% performance on contract satisfaction.
These findings motivate contract-aware evaluation on benchmarks with explicit contracts, where CVTs are executed rather than filtered out. Because this evaluation assumes that the provided contracts accurately reflect the task specification, we assess contract–specification alignment and filter out misaligned cases.
3.1 HumanEval+ and MBPP+ for Contract-Aware Analysis
HumanEval+ and MBPP+ annotate each task with an explicit contract field alongside the natural language description and reference solution. This design enables analyses that are difficult to perform in standard code generation benchmarks. First, we can directly test contract satisfaction by executing implementations on ill-formed inputs that violate the contract and checking whether they are rejected. Second, we can validate the contracts themselves by comparing their semantics with the task specification and determining whether they are specification-aligned. However, standard evaluation pipelines are not equipped to leverage these opportunities, as they are designed to use only well-formed inputs. By addressing this gap, we utilize HumanEval+ and MBPP+ as a natural foundation.
3.2 Contract–Specification Alignment Check
Contract satisfaction evaluation assumes that each contract captures mandatory input constraints stated or implied by the task description. If a contract is irrelevant, overly restrictive, or internally contradictory, contract-based evaluation becomes noisy. We therefore assess contract–specification alignment using an LLM judge and exclude tasks with misaligned contracts before downstream analyses.
Figure 3: Overview of ContractEval construction.
4 ContractEval Benchmark
We introduce ContractEval, a code-generation benchmark with contract-violating tests (CVTs) to evaluate whether generated program reject invalid inputs specified by assertion-level contracts. Figure 3 illustrates the detail of our neuro-symbolic construction pipeline using both an LLM and SMT solver. Each ContractEval task consists of: (i) the original benchmark prompt and well-formed functional tests for pass@k, (ii) contract-aware query reconstructions by injecting intended contract information to original queries. and (iii) verified CVTs derived from reference assertions for contract satisfaction. We further introduce new evaluation metrics to quantify contract satisfaction under contract-violating tests.
4.1 Contract-Violating Test Construction
CVT Synthesis Pipeline.
We propose a neuro-symbolic pipeline that combines the semantic translation capability of LLMs with the satisfiability guarantees of SMT solvers. The pipeline produces CVTs that violate a chosen subset of contracts while remaining feasible with respect to the remaining contracts within two stages. The first stage converts natural language level contract assertions into SMT-LIB constraints defined over an Algebraic Data Type (ADT). The second stage enumerates combinations of contracts, checks satisfiability for each combination, and synthesizes concrete inputs from satisfying combinations.
Formalization with a Canonical ADT Template.
Programs may accept inputs that combine primitive and structured values, including nested containers with mixed element types. Such inputs are difficult to encode directly in SMT because most solvers assume statically typed signatures and require explicit constructors for structured data. We address this challenge with a canonical value ADT embedded in an SMT-LIB template, as detailed in Appendix F. The ADT provides a unified representation for integers, floats, strings, booleans, and nested containers, allowing contracts to be encoded as SMT constraints over a single representation using shared constructors and auxiliary predicates. This uniform formalization supports systematic construction of constraint sets for different combinations of contract satisfaction and violation, which enables query generation for all non-empty violation combinations.
An LLM then populates the template by converting the reference contracts into ADT-level constraints and instantiating the corresponding template slots. The instantiated template specifies the contract constraints for the solver on top of the fixed ADT representation. This template-driven translation yields a syntactically valid SMT-LIB instance and preserves the semantics of contracts required for the solver to recognize.
Combinatorial Synthesis via SMT Solving.
The pipeline performs combinatorial synthesis to generate test cases that cover distinct and logically feasible violation scenarios implied by the contract set. For a task with $n$ contracts, we consider all $2^{n}-1$ non-empty subsets as violation targets. For each target subset, we construct an SMT query that negates the targeted contract predicates and enforces satisfaction of the remaining predicates. The solver checks satisfiability and discards UNSAT targets that conflict with the remaining constraints. For each SAT query, the solver returns a ADT-level model assignment, which constitutes a CVT in the encoded space. We decode this assignment into an executable test input by converting ADT constructors into their corresponding runtime values and recursively rebuilding nested containers. The resulting CVTs are precisely targeted and logically consistent.
Feasibility Filtering for CVTs.
Test cases synthesized by the SMT solver are logically consistent with respect to the contract predicates. Logical consistency alone does not ensure that an input is suitable for evaluating contract satisfaction. A valid CVT must be functionally feasible while still violating the intended contract. The same input should execute successfully when contract checks are removed, and it should fail with the intended assertion when those checks are enabled. Inputs that crash with generic runtime exceptions, such as TypeError or IndexError, even without contracts, do not isolate contract enforcement and instead reflect functional invalidity.
We filter synthesized CVTs by executing each input on the reference implementation under two configurations: (i) with assertions removed to verify functional feasibility, and (ii) with assertions intact to confirm the targeted contract violation. This filtering reduces the task set from 419 to 364 because some tasks admit no input that is both executable without contracts and contract violating with assertions enabled. The resulting tasks ensure that ContractEval measures contract enforcement rather than robustness to generic errors.
4.2 Contract-Aware Query Reconstruction
Each task in ContractEval includes an original natural-language query and a set of reference contract assertions. We use a commercial LLM to reconstruct a contract-aware query by interpreting the intent of each assertion and rewriting the original query so that the corresponding input constraints are stated explicitly in natural language. This step standardizes how contract information is presented to code generation models and supports controlled comparisons by ensuring that evaluations focus on the model’s ability to satisfy explicitly stated constraints.
4.3 ContractEval Metrics
ContractEval reports CVT quality metrics, assertion violation capture (AVC) and target specificity (TS) to validate that synthesized tests are clause-targeted rather than generic crashing inputs. ContractEval also reports functional correctness on well-formed inputs using pass@k, and contract satisfaction of generated programs using contract satisfaction rate (CSR) below.
CVT Quality Metrics (AVC, TS).